OpenStudy (anonymous):

State any restrictions on the variables. 9x-3x^2/x+3 A. x ≠ 3 B. x ≠ –3 C. x ≠ –3 and x ≠ 3 D. no restrictions are needed

OpenStudy (anonymous):

@davidmack

OpenStudy (tkhunny):

9x-3x^2/x+3 As written, \(9x - \dfrac{3x^{2}}{x} + 3\), we need \(x\ne 0\) If you remember your Order of Operations and use some parentheses ( 9x-3x^2)/(x+3) = \(\dfrac{9x-3x^{2}}{x+3}\), we need \(x \ne -3\) Proper Notation makes all the difference.

OpenStudy (anonymous):

x cant = 3 so a i think

OpenStudy (anonymous):

holy crap thx!

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

So was it x≠3 or x≠-3 or both?

OpenStudy (tkhunny):

Why would x = 3 be a problem?